# Diversity of rocky planet atmospheres in the C-H-O-N-S-Cl system with interior dissolution, non-ideality, and condensation: Application to TRAPPIST-1e and sub-Neptunes

### v1

@@ Appendix A Partially molten mantle
We repeat the analysis presented in Section 3.1 for a partially molten TRAPPIST-1e with a mantle melt fraction of 10%, whereby the influence of solubility is lessened compared to a fully molten mantle (compare Figure 1, 2, 4, 5 to Figure A1, A2, A3, A4, respectively). Even with just 10% melt fraction, C/H in the atmosphere is boosted and its distribution broadened compared to total C/H (Figure 3). Compared to a fully molten mantle, a partially molten mantle permits a greater diversity of atmospheres since a broader range of gas mixtures comprised of hydrogen and carbon species are possible. Although larger H2O atmospheres can occur, the median H2O in the atmosphere is still only 11 mol.% versus 1 mol.% for a fully molten mantle. The mean molar mass (Figure A1, bottom right) reflects this greater atmospheric diversity, where a subsidiary peak around 5 g/mol represents H2-dominated atmospheres at reduced conditions in a mixture with CH4 and CO. The CO peak at 28 g/mol is evident, albeit now supplemented by mixtures of H2O (18 g/mol) and CO2 (44 g/mol) in approximately equal proportion. The molar mass distribution also has an extended tail, driven by the prevalence of CO across a broad range of oxygen fugacity as well as the emergence of CO2 (44 g/mol) and SO2 (64 g/mol) as prevalent oxidized species. Regardless, the thermodynamic stability of CO over a range of O2 ensures that it remains a dominant species.
@@
Figure A4 shows the atmospheres at 280 K derived from the partially molten starting state. Compared to those derived from the fully molten starting state, the atmospheres exhibit the same complex relationships albeit the solution space is enlarged for certain quantities. This is because the smaller initial C/H allows H2 and CH4 atmospheres to be more prevalent at reduced starting conditions which can produce mixtures of hydrogen and carbon species. Nevertheless, CO2 atmospheres are dominant in similarity with atmospheres derived from the fully molten starting state. H2O is only present in the atmosphere below the percent level since water condenses to produce a median surficial reservoir of 0.4 Earth oceans by mass (EO) with a maximum inventory around 5.5 EO, which is almost a ten fold increase compared to the fully molten starting condition (Figure 5). -sulfur condensation can produce up to around 0.4 EO, although an expected amount is closer to 0.1 EO. In contrast, the total graphite reservoir is usually around 0.3 EO, but for reduced starting conditions can reach up to 10 EO. Large reservoirs of all three condensates can coexist, usually for starting conditions around the IW buffer.
The overall trends of atmospheres above planetary surfaces with stable condensates are comparable to the fully molten starting state in terms of the dominance of CO2 and CH4 atmospheres (compare Figure A5 and 9). However, there are a few notable differences. First, most water oceans are stable beneath CO2 atmospheres and most graphite reservoirs are stable beneath CH4 atmospheres. This trend is reversed compared to the atmospheres derived from a fully molten starting state. Second, H2S atmospheres are absent for the fully molten starting state, but now feature above all three condensate reservoirs. Third, SO2 atmospheres can support up to 3 EO of water as well as 0.15 EO of -sulfur. Finally, N2 atmospheres, although low in number, are also compatible with large reservoirs of water and -sulfur. Figure A6 shows the atmospheric speciation in equilibrium with stable reservoirs of water and graphite only, which excludes -sulfur and ammonium chloride compared to Figure 10 in the main text.
Tables A1 and A2 summarise the atmosphere speciation for TRAPPIST-1e at 280 K for atmospheres with stable water, graphite, and -sulfur, for fully molten and partially molten starting states, respectively (Figure 10).
@@
Appendix B Comparison with FactSage 8.2
We independently verified the results of our model by comparison to FactSage calculations (version 8.2, Bale et al., 2016), assuming ideal gas behaviour (Table B1). This comparison reveals that the partial pressure of a given volatile differs by at most a few percent. Additional comparison cases are available in the open source code including cases with condensates.
@@
3 | 1 | -2.0 | 1400 | 6.2 | 0.2 | 175 | 13.8 | 38.1 | 6.2 | 0.2 | 176 | 13.8 | 38 |
3 | 1 | +0.5 | 1400 | 46.6 | 30.7 | 237 | 337 | 28.7 | 46.4 | 30.9 | 237 | 337 | 28.7 |
1 | 0.1 | +2.0 | 1400 | 0.9 | 3.2 | 27.8 | 218 | 0 | 0.9 | 3.3 | 27.4 | 218 | 0 |
1 | 5 | +4.0 | 1400 | 9.6 | 358 | 5.4 | 432 | 0 | 10.21 | 357 | 5.8 | 432 | 0 |
1 | 1 | 0 | 873 | 0 | 0 | 59 | 18.3 | 19.5 | 0 | 0 | 59 | 18.3 | 19.5 |
Appendix C Bounded real gas equation of state
@@
Equation C2 can be included as part of a volume integral calculation when , from which the fugacity can be determined using
Appendix D Treatment of oxygen
The condensed reservoir (molten/solid rock) of a rocky planet dominates its gaseous reservoir in terms of mass or number of moles. During the magma ocean stage, in which the molten mantle and atmosphere communicate due to similar dynamic timescales, it is therefore expected that mantle chemistry plays a role in regulating the chemistry of the combined atmosphere-mantle system. In particular, oxygen fugacity (O2) is a crucial parameter in mantle geochemical systems because it governs the oxidation state of elements, influencing mineral stability, phase equilibria, and the speciation of volatiles such as hydrogen, carbon, and sulfur. It is reasonable to assume that oxygen fugacity (O2) is buffered by the iron-wüstite (IW) equilibrium in the early history of a rocky planet because planetary accretion and differentiation involve metal-silicate equilibration under reducing conditions. During core formation, metallic iron segregates from the silicate mantle, and the prevailing redox state is largely controlled by Fe-FeO equilibrium, which defines the IW buffer. Rock analyses confirm the IW buffer as regulating the redox state of Earth’s magma ocean (Sossi et al., 2020).
Hence for the high temperature calculations of both TRAPPIST-1e and K2-18b the control of mantle chemistry on atmospheric speciation is encapsulated in the constraint on O2, which is defined relative to the IW buffer that exclusively determines (absolute) O2 (in bar) at a given temperature and a reference pressure of 1 bar. Representing O2 relative to the IW buffer does not necessitate that the atmosphere of K2-18b is actually buffered by the Fe–FeO reaction at present day. Furthermore, it should be appreciated that imposing a buffered oxygen constraint does not in itself constrain the total oxygen abundance. This is evident, for example, by considering the following redox reaction:
@@ Appendix D Treatment of oxygen
where is temperature and fugacity, in which the later can be considered equivalent to partial pressure for this discussion. If O2 is constrained relative to the IW buffer then both O2 and are solely functions of temperature and these collectively govern the ratio of H2 to H2O but importantly not their absolute values. Imposing an additional hydrogen abundance constraint (or alternatively pressure constraint) is required to break the degeneracy to allow a unique determination of H2 and H2O. That is, 2 equations (Equation D2 and hydrogen mass balance) can be solved to determine 2 unknowns (H2 and H2O), implicitly constraining the total oxygen abundance. Hence if O2 is prescribed as a constraint, Atmodeller solves for the total oxygen abundance self-consistently as part of the solution process. For the cooled atmosphere calculations of TRAPPIST-1e, the elemental abundance of species in the atmosphere (including oxygen) are prescribed as constraints, in which case Atmodeller then ensures a self-consistent determination of O2.
Appendix E Interplay of solubility and non-ideality
@@
where is total pressure, is a constant that regulates non-ideality and bar is the pressure at which the volatile behaves ideally. Hence ideal behaviour is recovered as or . For simplicity, temperature-dependence can also be accounted for through the choice of P0. Substituting Equations E2 and E3 into E1 gives:
@@
and is the mass of the rocky planet, i.e. silicate mantle plus iron core. Equation E4 can be simplified for a single species atmosphere () and for , appropriate for H2 (Hirschmann et al., 2012):
@@
Appendix F Additional K2-18b Figures
Figures F1 and F2 show how atmospheric pressures (a), atmospheric C/H and C/O ratios by mass (b), solubilities (c) and fugacity coefficients (d) vary with oxygen fugacity and total C/H mass ratio, respectively, between the real and ideal cases. In these simulations, the planet mass, surface radius and temperature are the same as those in Section 3.2 and Figures 6 and 7, and only one input parameter is varied (either oxygen fugacity or C/H mass ratio) while the others are kept constant.

### v2

@@ Appendix B Partially molten mantle
We repeat the analysis presented in Section 3.1 for a partially molten TRAPPIST-1e with a mantle melt fraction of 10%, whereby the influence of solubility is lessened compared to a fully molten mantle (compare Figure 1, 2, 4, 5 to Figure B1, B2, B3, B4, respectively). Even with just 10% melt fraction, C/H in the atmosphere is boosted and its distribution broadened compared to total C/H (Figure 3). Compared to a fully molten mantle, a partially molten mantle permits a greater diversity of atmospheres since a broader range of gas mixtures comprised of hydrogen and carbon species are possible. Although larger H2O atmospheres can occur, the median H2O in the atmosphere is still only 11 mol.% versus 1 mol.% for a fully molten mantle. The mean molar mass (Figure B1, bottom right) reflects this greater atmospheric diversity, where a subsidiary peak around 5 g/mol represents H2-dominated atmospheres at reduced conditions in a mixture with CH4 and CO. The CO peak at 28 g/mol is evident, albeit now supplemented by mixtures of H2O (18 g/mol) and CO2 (44 g/mol) in approximately equal proportion. The molar mass distribution also has an extended tail, driven by the prevalence of CO across a broad range of oxygen fugacity as well as the emergence of CO2 (44 g/mol) and SO2 (64 g/mol) as prevalent oxidized species. Regardless, the thermodynamic stability of CO over a range of O2 ensures that it remains a dominant species.
@@
Figure B4 shows the atmospheres at 280 K derived from the partially molten starting state. Compared to those derived from the fully molten starting state, the atmospheres exhibit the same complex relationships albeit the solution space is enlarged for certain quantities. This is because the smaller initial C/H allows H2 and CH4 atmospheres to be more prevalent at reduced starting conditions which can produce mixtures of hydrogen and carbon species. Nevertheless, CO2 atmospheres are dominant in similarity with atmospheres derived from the fully molten starting state. H2O is only present in the atmosphere below the percent level since water condenses to produce a median surficial reservoir of 0.4 Earth oceans by mass (EO) with a maximum inventory around 5.5 EO, which is almost a ten fold increase compared to the fully molten starting condition (Figure 5). -sulfur condensation can produce up to around 0.4 EO, although an expected amount is closer to 0.1 EO. In contrast, the total graphite reservoir is usually around 0.3 EO, but for reduced starting conditions can reach up to 10 EO. Large reservoirs of all three condensates can coexist, usually for starting conditions around the IW buffer.
The overall trends of atmospheres above planetary surfaces with stable condensates are comparable to the fully molten starting state in terms of the dominance of CO2 and CH4 atmospheres (compare Figure B5 and 9). However, there are a few notable differences. First, most water oceans are stable beneath CO2 atmospheres and most graphite reservoirs are stable beneath CH4 atmospheres. This trend is reversed compared to the atmospheres derived from a fully molten starting state. Second, H2S atmospheres are absent for the fully molten starting state, but now feature above all three condensate reservoirs. Third, SO2 atmospheres can support up to 3 EO of water as well as 0.15 EO of -sulfur. Finally, N2 atmospheres, although low in number, are also compatible with large reservoirs of water and -sulfur. Figure B6 shows the atmospheric speciation in equilibrium with stable reservoirs of water and graphite only, which excludes -sulfur and ammonium chloride compared to Figure 10 .
Tables B1 and B2 summarise the atmosphere speciation for TRAPPIST-1e at 280 K for atmospheres with stable water, graphite, and -sulfur, for fully molten and partially molten starting states, respectively (Figure 10).
@@
Appendix C Comparison with FactSage 8.2
We independently verified the results of our model by comparison to FactSage calculations (version 8.2, Bale et al., 2016), assuming ideal gas behaviour (Table C1). This comparison reveals that the partial pressure of a given volatile differs by at most a few percent. Additional comparison cases are available in the open source code including cases with condensates.
@@
| 3 | 1 | -2.0 | 1400 | 6.2 | 0.2 | 175 | 13.8 | 38.1 | 6.2 | 0.2 | 176 | 13.8 | 38 |
| 3 | 1 | +0.5 | 1400 | 46.6 | 30.7 | 237 | 337 | 28.7 | 46.4 | 30.9 | 237 | 337 | 28.7 |
| 1 | 0.1 | +2.0 | 1400 | 0.9 | 3.2 | 27.8 | 218 | 0 | 0.9 | 3.3 | 27.4 | 218 | 0 |
| 1 | 5 | +4.0 | 1400 | 9.6 | 358 | 5.4 | 432 | 0 | 10.21 | 357 | 5.8 | 432 | 0 |
| 1 | 1 | 0 | 873 | 0 | 0 | 59 | 18.3 | 19.5 | 0 | 0 | 59 | 18.3 | 19.5 |
Appendix D Bounded real gas equation of state
@@
Equation D2 can be included as part of a volume integral calculation when , from which the fugacity can be determined using
Appendix E Treatment of oxygen
The condensed reservoir (molten/solid rock) of a rocky planet dominates its gaseous reservoir in terms of mass or number of moles. During the magma ocean stage, in which the molten mantle and atmosphere communicate due to similar dynamic timescales, it is therefore expected that mantle chemistry plays a role in regulating the chemistry of the combined atmosphere-mantle system. In particular, oxygen fugacity (O2) is a crucial parameter in mantle geochemical systems because it governs the oxidation state of elements, influencing mineral stability, phase equilibria, and the speciation of volatiles such as hydrogen, carbon, and sulfur. It is reasonable to assume that oxygen fugacity (O2) is buffered by the iron-wüstite (IW) equilibrium in the early history of a rocky planet because planetary accretion and differentiation involve metal-silicate equilibration under reducing conditions.
Rock analyses confirm the redox state of Earth’s magma ocean was close to the IW buffer (Sossi et al., 2020). Hence for the high temperature calculations of both TRAPPIST-1e and K2-18b the control of mantle chemistry on atmospheric speciation is encapsulated in the constraint on O2, which is defined relative to the IW buffer that exclusively determines (absolute) O2 (in bar) at a given temperature and a reference pressure of 1 bar. Representing O2 relative to the IW buffer does not necessitate that the atmosphere of K2-18b is actually buffered by the Fe–FeO reaction at present day. Furthermore, it should be appreciated that imposing a buffered oxygen constraint does not in itself constrain the total oxygen abundance. This is evident, for example, by considering the following redox reaction:
@@ Appendix E Treatment of oxygen
where is temperature and fugacity, in which the later can be considered equivalent to partial pressure for this discussion. If O2 is constrained relative to the IW buffer then both O2 and are solely functions of temperature and these collectively govern the ratio of H2 to H2O but importantly not their absolute values. Imposing an additional hydrogen abundance constraint (or alternatively pressure constraint) is required to break the degeneracy to allow a unique determination of H2 and H2O. That is, 2 equations (Equation E3 and hydrogen mass balance) can be solved to determine 2 unknowns (H2 and H2O), implicitly constraining the total oxygen abundance. Hence if O2 is prescribed as a constraint, Atmodeller solves for the total oxygen abundance self-consistently as part of the solution process. For the cooled atmosphere calculations of TRAPPIST-1e, the elemental abundance of species in the atmosphere (including oxygen) are prescribed as constraints, in which case Atmodeller then ensures a self-consistent determination of O2.
Appendix F Interplay of solubility and non-ideality
@@
where is total pressure, is a constant that regulates non-ideality and bar is the pressure at which the volatile behaves ideally. Hence ideal behaviour is recovered as or . For simplicity, temperature-dependence can also be accounted for through the choice of P0. Substituting Equations F2 and F3 into F1 gives:
@@
and is the mass of the rocky planet, i.e. silicate mantle plus iron core. Equation F4 can be simplified for a single species atmosphere () and for , appropriate for H2 (Hirschmann et al., 2012):
@@
Appendix G Additional K2-18b Figures
Figures G1 and G2 show how atmospheric pressures (a), atmospheric C/H and C/O ratios by mass (b), solubilities (c) and fugacity coefficients (d) vary with oxygen fugacity and total C/H mass ratio, respectively, between the real and ideal cases. In these simulations, the planet mass, surface radius and temperature are the same as those in Section 3.2 and Figures 6 and 7, and only one input parameter is varied (either oxygen fugacity or C/H mass ratio) while the others are kept constant.